Suppose Ginny gets a sales bonus at her place of work that gives her an extra $600 of disposable income. She chooses to spend $360 and save the remaining $240.

From this, you can tell that Ginnys marginal propensity to consume (MPC) is ______ and her marginal propensity to save is ____

Mathematically, it must always be true that:

Saving = Disposable Income + Consumption

Consumption - Disposable Income

Disposable Income - Consumption

Therefore, it must also be true that:

MPS = 1 + MPC

1 - MPC

MPC
Answer #1

MPC = change in consumption/change in income= 360/600 = 0.6

MPS = change in saving/change in income or 1-MPC = 1-0.6=0.4

Mathematically, Disposable income = Consumption+savings so, 600 = 360+240

saving = disposable income - consumption

Therefore, MPS = 1-MPC
Answer #2
0.80 0.20 disposable - savings 1 + MPS
source: math
answered by: jj